Solve using substitution method
7x+2y=-30
x=42-8y

Answer by mgmoeab(37)

You are given:
(1) 7x+2y=-30
(2) x=42-8y
Substitute (2) in (1)
7 ( 42 -8y) + 2y = -30
Distributing:
294 - 56y + 2y = -30
Combining like-terms:
294 - 54y = -30
-54y = -324
y = 6